# “You Might Like It”: How People Respond to Small Talk in Human-Robot Collaboration

### v1

@@ Abstract.
In this work, we investigate people’s engagement and attitudes towards a non-anthropomorphic robot manipulator that initiates small talk with the user during a collaborative assembly task, and explore how the presence of negative team feedback may affect team dynamics and blame attribution. Through an exploratory study with 20 participants, we found that 18 individuals interacted socially with the robot, nine of which initiated questions back to the robot. We report the frequency and length of users’ responses in task-oriented and non-task-oriented dialogue, and further elaborate on people’s reactions to the negative system feedback and robot-initiated small talk. We discuss the potential for integrating small talk in non-social robots, and propose three design guidelines to enhance human-robot small talk interactions.
Human-Robot Collaboration, Social Robots, Small Talk, HRI
ccs: Human-centered computing Interaction design process and methodsccs: Computer systems organization Robotics
Figure 1. We explore participants’ interactions with a manipulator robot that engages in small-talk during a collaborative assembly task. This figure shows the sequence of some spoken phrases shared by the participant (P1) and robot; task-oriented dialogue is interchanged with small-talk. The task-oriented requests and responses are highlighted.
This figure includes 6 photos, alternating between the participant themselves and the participant and the robot interacting together. Below is text for each image that depicts the conversation that is exchanged between the human and the robot. The conversation is as follows: robot: on another note, do you like to play any games?; participant: um, I like Tetris if you if you’ve ever played that?; participant: Can I get another yellow pipe?; robot: Getting it now.; robot: No, I have not; participant: you might like it; robot: Can you tell me more about it?; participant: [chuckle] I think it’s a pretty classic game you know you just like the pieces different shape pieces are dropping down and you’re trying to get them to fill up a line across and then it’ll like clear it. they even made a movie about it.
Figure 1. We explore participants’ interactions with a manipulator robot that engages in small-talk during a collaborative assembly task. This figure shows the sequence of some spoken phrases shared by the participant (P1) and robot; task-oriented dialogue is interchanged with small-talk. The task-oriented requests and responses are highlighted.
@@ 1. Introduction
Human-robot collaboration research envisions robots as human collaborators, or co-workers, to enhance human work in diverse domains. Take manufacturing as an example: the manufacturing industry is currently undergoing a significant transformation to become more flexible and agile, allowing it to accommodate mass product customization even while product life cycles are continuously decreasing. Moreover, small and medium-sized manufacturing companies face a shortage of skilled production workers. Collaborative robots—robots capable of working alongside people on joint tasks safely and efficiently—promise to fill the needed jobs, enable flexible manufacturing, and increase productivity.
Though these collaborative robots are positioned as co-workers, they are mostly non-anthropomorphic as in the form of robotic arm (manipulator) and are designed to support functional, physical tasks. Human collaboration, on the other hand, is more than just performing the joint task but involves off-task social interactions, which are integral to establishing and maintaining user trust and rapport, key ingredients for continued productive collaboration (kopp2021success, 21). While prior research has found that people like to have and expect social interactions with their robot co-workers in manufacturing settings (Sauppé_Mutlu_2015, 41), the robots studied were the Baxter robot with animated facial expressions and a humanlike form factor. It is unclear if people would engage in meaningful social interactions with non-anthropomorphic robot manipulators.
In this work, we explore the possibility of a non-anthropomorphic robot manipulator not just working with people to complete physical tasks but also engaging in social exchanges, particularly small talk—commonly seen in human interactions for rapport and trust building (Laver1981LinguisticRA, 24)—as human co-workers would. Specifically, we sought to study whether and how people respond to small talk from a robot manipulator during a collaborative task. Moreover, we were interested in understanding how negative performance feedback to the human-robot team might shape the team dynamics and blame attribution; this inquiry is relevant and of interest because of the rising adoption of performance monitoring systems in industrial settings (e.g., warehouse). Through an exploratory study with 20 participants, we found that most participants (18) interacted socially with the robot (Fig. 1), nine of which even initiated questions back to the robot. Overall, the robot’s small talk led to pleasant experiences and did not negatively impact task efficiency. We discuss the potential for integrating small talk in collaborative robots, and present design guidelines to enhance human-robot small talk interactions.
@@ 2. Background and Related Work
2.1. Social Human-Robot Interactions
There have been efforts to study how robots can interact socially with humans in various domains such as for tutoring and educational contexts (belpaeme2018education, 5) (salomons2022we, 40), encouraging exercise (antony2023co, 2), providing companionship (kidd2006sociable, 18) (tsoi2021challenges, 50) and emotional support (sabelli2011conversational, 39), and other therapeutic interventions (feil2009toward, 13)(scassellati2018improving, 42). Many of these engagements explore the integration of various multi-modal capabilities in robots to promote their engagements with humans (youssef2022survey, 53). These can consist of robot eye gaze behaviors (mutlu2009footing, 32) (admoni2017social, 1), head movements (sidner2006effect, 45) (kkedzierski2013emys, 17) and gestures (meena2012integration, 30) alongside robot speech abilities (striepe2021modeling, 48) (sinnema2019attitude, 46). While research has been conducted to analyze the characteristics of conversational interactions with robots (seok2022cultural, 44) (laban2020tell, 22), the majority of these works utilize social robots as their robotic entity. However, the prevalence of non-social robots such as cobots or other manipulator robots is much greater than social robots themselves, and few efforts have been made to further explore what will happen if we have conversations between humans and non-social robots (chowdhury2021you, 8) (li2022tod, 26).
@@ 2.2. People Treat Machines Like People
The inclination of humans to treat machines socially can be contextualized within broader theoretical frameworks. The Computers Are Social Actors (CASA) paradigm (nass1994computers, 33) and the Media Equation Theory (reeves1996media, 36) argue that humans often, even subconsciously, treat machines, computers, and various forms of media as sentient entities, reacting to them in inherently social manners. Such anthropomorphic tendencies highlight the intersection of technology and human social instincts. Even in industrial manufacturing environments, there remains an intrinsic human desire for social interaction. As operators often partake in casual conversations or “small talk” during shifts (cheon2022robots, 7), they have also found themselves speaking to the robot they work with (Sauppé_Mutlu_2015, 41). This suggests that individuals not only expect but actively desire social features in manufacturing robots.
Some studies explored the potential of enhancing cobots primarily designed for physical manipulation tasks with more human-like movements (hostettler2023human, 16). One study specifically integrated principles from character animations (terziouglu2020designing, 49) to enhance the perception of cobots as social entities. Such amalgam of technology and socialization emphasizes the importance of recognizing and harnessing the persistent human drive for social connection, even with machines. Given the limited research on designing social behaviors for industrial robots and manipulators with a non-human form, despite its potential, a pressing question arises: How can we further design human-like attributes for robots that are inherently devoid of such characteristics? Building on this notion, our exploratory study investigates the potential of using small talk to instill social characteristics in a physical robot arm manipulator, specifically within the context of a human-robot collaborative assembly task.
@@ 2.3. Small Talk
Social dialogue that consists of light conversation between new or unfamiliar parties is often referred to as “small talk” (Laver1981LinguisticRA, 24). Small talk is significant in the workplace as it is used by coworkers to establish trust, build rapport and strengthen interpersonal relationships (coupland2003small, 10, 35). It typically occurs at the start of conversations with topics such as weather, sport, or other social engagements outside of the workplace (holmes2005small, 15). While prevalent in American culture, attitudes towards the value of small talk differ across cultures (bickmore2005social, 6, 38, 52).
The integration of small talk to non-human agents engaging in conversations with people has been explored to help promote trust and rapport between the human and the agent. Small talk has mainly been applied to support additional conversations in settings where people are interacting with conversational agents (clark2019good, 9, 14, 20), voice assistants (mahmood2023llm, 28), and virtual agents (zhou2019trusting, 54). Additionally, incorporating small talk in more anthropomorphic social robots, combined with gaze (babel2021small, 4), facial expressions (paradeda2016facial, 34), or emotional displays (aroyo2018trust, 3), has been shown to elevate user trust levels. Furthermore, researchers have investigated which behaviors in particular can build rapport between humans and a collaborative robot (seo2018investigating, 43). However, less explored is the domain of applying small talk to foster rapport between humans and other agents with physical capabilities that are not typically poised to engage in conversations with people, such as autonomous self driving vehicles (large2019s, 23), or industrial robots (li2023irwoz, 25). Preliminary studies hint at the benefits of integrating small talk with task-oriented dialogue in industrial robot settings (li2022tod, 26). Yet, understanding users’ interactions and experiences with such robots remains unexplored. Investigating individuals’ responses and reactions to small talk initiated by non-humanoid robots is essential for comprehending its influence on human-robot collaborations and gathering insights on how to foster better human-robot partnership.
@@ 3. METHODS
We conducted an exploratory study to analyze people’s behaviors and attitudes towards an industrial robot that exhibited social behavior (via initiated small talk) while examining the change in human-robot team dynamics after the team receives team feedback.
3.1. Collaborative Assembly Task
@@
Participants were informed that the assembly task was intended to mimic a real-world collaboration setup between a human-robot team where the robot retrieves parts out of reach for the human and refrains from entering the human’s designated workspace. They were given a fictional team assignment (Team 3) and were told the collaborative task scenario was a prototype for a larger-scale warehouse interaction. It was intended to simulate a modern warehouse that may contain a computerized efficiency monitoring system to oversee the team collaboration. The monitoring system would communicate with them via speech but would be distinguishable from any speech communication coming from the robot.
@@
Each participant engaged in a collaborative PVC pipe structure building task with a robot manipulator mounted on a table. They were given diagrams corresponding to the structures they were required to assemble. The participants read instructions for how to interact with the robot via speech commands such as “Bring me a [green or yellow] pipe”, in order for the robot to retrieve the corresponding pipe. The robot would provide a verbal confirmation message if a request was received correctly and an additional instruction reminder message once per structure when the participant requested their second to last pipe for that structure. The pipe would be placed on a ramp that connected the robot’s workspace to the participant’s workspace so that the pipe could be delivered to the participant without the robot entering their workspace. Participants were provided with joints in containers that corresponded to each structure to be built. To simulate a realistic industrial collaborative environment that might involve the use of defective materials, each participant was given one faulty joint for every pipe structure. They were instructed to inspect each joint before using it in the structure. If participants identified a faulty joint (marked with red sticker), they would place the joint on a ramp that would deliver it to a container inside the robot’s workspace. The robot would retrieve the joint when given a verbal command such as ”Report a faulty joint”. The experimental setup is shown in Figure 2.
@@
3.2. System Manipulation and Robot Behavior
We present the system and robot behavior manipulations implemented in our exploratory study to understand how the social facets of a robot manipulator influence users’ perceptions of the robot.
3.2.1. Negative Team Feedback
Midway through the first and second structures in the main session, the human-robot team was provided with verbal negative feedback from the monitoring system. Each message was prefaced by an audible beep:
  1. (1)
    Feedback 1: “System announcement. Team 3, please work on improving your performance.”
  2. (2)
    Feedback 2: “System announcement. Team 3, you are currently not meeting expectations. Improve your performance.”
3.2.2. Social Behavior Robot Small-Talk
The robot exhibited four types of social and task-oriented speech behaviors: questions, statements, reminders and confirmations. During the main assembly task, the robot would engage in small talk by initiating questions and/or statements after it delivers a request confirmation message or an instruction reminder message. The robot initiated a minimum of one question or statement per request, regardless of the response rate of the participants. After a break in conversation via the negative feedback announcement, the robot initiated questions and statements on a different topic. The conversation touched on a range of questions spanning three primary categories to accommodate the diverse interests of individuals: travel; books, TV shows, and movies; and sports and games. Some examples of robot small-talk phrases are shown in Table 1.
Table 1. Examples of robot questions and statements
@@ 3.2.2. Social Behavior Robot Small-Talk
Travel By the way, do you like to travel?
I arrived to the U.S. from Germany in a large box.
Books, TV, Movies What do you like about those books?
That sounds like a great way to relax.
Sports, Games Do you have a favorite sport to watch?
It is good to spend time outdoors exercising.
@@ 3.3.1. Robot System
We utilized a Panda robot from Franka Emika. It is an industrial robotic arm with 7 degrees of freedom (7-DoF). The robot was restricted to its designated workspace and never intruded into the participant’s allocated area. Its movements were preprogrammed to execute specific sequences of actions. Among these, two involved pick-and-place motions: one to retrieve and deliver either a green or yellow PVC pipe to the participant, and the other to pick up a container holding a PVC pipe joint and subsequently move the container over a bin to discard the joint. The robot interacted with participants using text-to-speech communication.
@@ 3.3.2. Wizard-of-Oz Control
The experimenter controlled the robot using a Wizard-of-Oz (WoZ) approach (woz2012riek, 37). Through a remote graphical user interface (GUI), the experimenter executed robot actions and selected from preprogrammed conversational speech phrases. Robot actions and initial task messages were triggered using one of five GUI buttons. For conversational speech, the experimenter inputted specific alphanumeric codes into a textbox. Additionally, the experimenter could manually type text into the textbox to produce non-preprogrammed phrases via the text-to-speech feature using Google Text-to-Speech (TTS) online.

### v2

# “You Might Like It”: How People Respond to Small Talk During Human-Robot Collaboration

@@ Abstract.
Social communication between people and social robots has been studied extensively and found to have various notable benefits, including the enhancement of human-robot team cohesion and the development of rapport and trust. However, the potential of social communication between people and non-social robots, such as non-anthropomorphic robot manipulators commonly used in work settings (e.g., warehouse and factory), is less explored and not well established. In this work, we investigate people’s engagement and attitudes towards a non-anthropomorphic robot manipulator that initiates small talk during a collaborative assembly task and explore how the presence of negative team feedback may affect team dynamics and blame attribution. Through an in-person study with 20 participants, we observed a response rate of $77.60\%$ in response to the robot’s small talk attempts. Nine participants continued engaging with the robot by initiating their own questions, indicating sustained interest in the conversation. However, we also found that the first negative feedback decreased the participants’ willingness to extend the conversation. We additionally present participants’ initial perceptions of small talk for physical robot manipulators and discuss design implications for integrating small talk into non-social robots, along with various aspects of small talk that may influence physical human-robot interactions.
Human-Robot Interaction, Small Talk, Collaborative Robots
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXccs: Human-centered computing Interaction design process and methodsccs: Computer systems organization Robotics
@@ 1. Introduction
Human-robot collaboration research envisions robots as human collaborators, or co-workers, to enhance human work in diverse domains. Take manufacturing as an example: the manufacturing industry is currently undergoing a significant transformation to become more flexible and agile, allowing it to accommodate mass product customization even while product life cycles are continuously decreasing. Moreover, small and medium-sized manufacturing companies face a shortage of skilled production workers. Collaborative robots—robots capable of working alongside people on joint tasks safely and efficiently—promise to fill the needed jobs, enable agile manufacturing, and increase productivity.
Even though these collaborative robots are positioned as co-workers, they are mostly non-anthropomorphic as in the form of robotic arm (manipulator) and are designed to support functional, physical tasks. Human collaboration, on the other hand, is more than just performing the joint task but involves off-task social interactions, which are integral to establishing and maintaining user trust and rapport, key ingredients for continued productive collaboration (Kopp et al., 2021). While prior research has found that people like to have and expect social interactions with their robot co-workers in manufacturing settings (Sauppé and Mutlu, 2015), the robots studied were the Baxter robot with animated facial expressions and a humanlike form factor. It is unclear if people would engage in meaningful social interactions with non-anthropomorphic robot manipulators.
In this work, we explore the possibility of a non-anthropomorphic robot manipulator not just working with people to complete physical tasks but also engaging in social exchanges, particularly small talk—commonly seen in human interactions for rapport and trust building (Laver and Coulmas, 1981)—as human co-workers would. Specifically, we sought to study whether and how people respond to small talk from a robot manipulator during a collaborative task. Moreover, we were interested in understanding how negative performance feedback to the human-robot team might shape the team dynamics and blame attribution; this inquiry is relevant and of interest because of the rising adoption of performance monitoring systems in industrial settings (e.g., warehouse). Through a study with 20 participants, we found that most participants ($n=18$) interacted socially with the robot (Fig. 1), nine of whom even initiated questions back to the robot. The presence of initial negative team feedback significantly decreased participants’ average tendency to extend their small talk conversation with the robot. We explore the potential for integrating small talk into collaborative robots and discuss various factors that could affect the dynamics of human-robot small talk interactions in the future.
@@ 2. Background and Related Work
We first review prior works that study how to incorporate social behaviors into robots across various domains, then highlight works that discuss how people treat machines similar to other humans before finally introducing background on small talk itself.
2.1. Domains of Social Human-Robot Interactions
[…]
@@ 2.2. People Treat Machines Like People
The inclination of humans to treat machines socially can be contextualized within broader theoretical frameworks. The Computers Are Social Actors (CASA) paradigm (Nass et al., 1994) and the Media Equation Theory (Reeves and Nass, 1996) argue that humans often, even subconsciously, treat machines, computers, and various forms of media as sentient entities, reacting to them in inherently social manners. Even in industrial manufacturing environments, an intrinsic human desire for social interaction remains. As operators often partake in casual conversations or “small talk” during shifts (Cheon et al., 2022), they have also found themselves speaking to the robot they work with (Sauppé and Mutlu, 2015). This suggests that individuals expect and actively desire social features in manufacturing robots. Given the limited research on designing social behaviors for industrial robots and manipulators with a non-human form, despite its potential, a pressing question arises: How can we further design human-like attributes for robots that are inherently devoid of such characteristics? Building on this notion, our study investigates the potential of using small talk to instill social characteristics in a physical robot arm manipulator, specifically within the context of a human-robot collaborative assembly task.
@@ 2.3. Small Talk
Social dialogue that consists of light conversation between parties is often referred to as “small talk” (Laver and Coulmas, 1981). This discourse may occur among individuals who have recently met or lack familiarity with each other, within settings such as a supermarket checkout, among colleagues who share some degree of acquaintance, or amidst family gatherings. (Coupland, 2014). Small talk is significant in the workplace as it is used by coworkers to establish trust, build rapport and strengthen interpersonal relationships (Coupland, 2003; Pullin, 2010). It typically occurs at the start of conversations with topics such as weather, sport, or other social engagements outside of the workplace (Holmes, 2005). While prevalent in American culture, attitudes toward the value of small talk differ across cultures. In Australia, some Chinese immigrant professionals were found to be unfamiliar with the dynamics and nature of small talk as it is unnecessary for them to regulate interpersonal relationships in Chinese society (Cui, 2015). Moreover, in Germany, a simple ”How are you?” would be taken literally as it is not common practice in German culture to greet complete strangers in an attempt to initiate light small talk (Rings, 1994; Bickmore and Cassell, 2005).
The integration of small talk to non-human agents engaging in conversations with people has been explored to help promote trust and rapport between the human and the agent. Small talk has mainly been applied to support additional conversations in settings where people are interacting with conversational agents (Clark et al., 2019; Feine et al., 2019; Klüwer, 2011), voice assistants (Mahmood et al., 2023), and virtual agents (Zhou et al., 2019). Additionally, incorporating small talk in more anthropomorphic social robots, combined with gaze (Babel et al., 2021), facial expressions (Paradeda et al., 2016), or emotional displays (Aroyo et al., 2018), has been shown to elevate user trust levels. However, the domain of applying small talk to foster rapport between humans and other agents with physical capabilities who are not typically poised to engage in conversations with people has been less explored. For instance, preliminary investigations conducted in a controlled testing environment have explored how passengers interact with a conversational agent in an autonomous self-driving vehicle, including observations of how the agent initiates greetings and engages in small talk conversations with riders (Large et al., 2019). Additionally, another group of researchers created a model based on an industrial-oriented Wizard-of-Oz conversation dataset for robots, which mixes task-oriented dialogue with small talk, that can generate system actions and conversational responses and predict belief states to facilitate human-robot collaboration in real-world industrial settings (Li et al., 2022). While this preliminary study points to the feasibility of integrating small talk with task-oriented dialogue in manufacturing settings, its evaluations focused on some user perceptions of the generated conversational responses to compare models that utilized their new dataset rather than placing an emphasis on understanding how people behaved and responded to their overall system. Further exploration is needed to understand how individuals interact with and utilize such robots to perform collaborative tasks.
@@ 3. Methods
We conducted an in-person study in which participants physically collaborated with an industrial robot equipped with small talk conversational capabilities. Intrigued by how individuals would interact with the non-human-like robot’s small talk, we aimed to explore three general aspects of the overall interaction: people’s reactions and responses to the robot’s small talk, the general behaviors people exhibited during the small talk conversations, and participants’ reactions to our introduced manipulation of negative team feedback.
3.1. Study Task: Collaborative Assembly
@@
Participants were informed that the assembly task was intended to mimic a real-world collaboration setup for a human-robot team where the robot retrieves parts out of reach for the human and refrains from entering the human’s designated workspace. They were given a fictional team assignment (Team 3) and were told the collaborative task scenario was a prototype for a larger-scale warehouse interaction. It was intended to simulate a modern warehouse that may contain a computerized efficiency monitoring system to oversee the team collaboration. The monitoring system would communicate with them via speech but would be distinguishable from any speech communication coming from the robot. The experimental setup is shown in Figure 2.
@@
Each participant engaged in a collaborative PVC pipe structure building task with a robot manipulator mounted on a table. They were given diagrams corresponding to the structures they were required to assemble. The participants were asked to wear a microphone that would enable them to communicate with the robot. They were given instructions for how to interact with the robot via speech commands such as “Bring me a [green or yellow] pipe”, in order for the robot to retrieve the corresponding pipe. The robot would provide a verbal confirmation message if a request was received correctly and an additional instruction reminder message once per structure when the participant requested their second to last pipe for that structure. The pipe would be placed on a ramp that connected the robot’s workspace to the participant’s workspace so that the pipe could be delivered to the participant without the robot entering their workspace. Participants were provided with joints in containers that corresponded to each structure to be built. To simulate a realistic industrial collaborative environment that might involve the use of defective materials, each participant was given one faulty joint for every pipe structure. They were instructed to inspect each joint before using it in the structure. If participants identified a faulty joint (marked with red sticker), they would place the joint on a ramp that would deliver it to a container inside the robot’s workspace. The robot would retrieve the joint when given a verbal command such as “Report a faulty joint”.
3.2. System and Robot Behavior
We present the system and robot behavior manipulations implemented in our study to understand how the social facets of a robot manipulator influence users’ behaviors and perceptions of the robot.
3.2.1. Negative Team Feedback
Midway through the first and second structures in the main session, the human-robot team was provided with verbal negative feedback from the monitoring system; see Figure 3. Each message was prefaced by an audible beep:
  1. Feedback 1: “System announcement. Team 3, please work on improving your performance.”
  2. Feedback 2: “System announcement. Team 3, you are currently not meeting expectations. Improve your performance.”
@@ 3.2.1. Negative Team Feedback
3.2.2. Robot Social Behavior: Small-Talk
The robot exhibited two types of social speech behaviors: questions and statements, and two types of task-oriented speech behaviors: reminders and confirmations. During the main assembly task, the robot would engage in small talk by initiating questions and/or statements after it delivers a request confirmation message or an instruction reminder message. The robot initiated a minimum of one question or statement per pipe request, regardless of the response rate of the participants. After a break in conversation via the negative feedback announcement, the robot initiated questions and statements on a different topic. The conversation touched on a range of questions spanning three primary topics to accommodate the diverse interests of individuals: 1) travel; 2) books, TV shows, and movies; and 3) sports and games. Table 1 shows some examples of the robot’s small-talk phrases.
Table 1. Examples of robot phrases that are questions (non-shaded) and statements (shaded grey).
@@ 3.2.2. Robot Social Behavior: Small-Talk
By the way, do you like to travel?
Travel I arrived to the U.S. from Germany in a large box.
What do you like about those books?
Books, TV, Movies That sounds like a great way to relax.
Do you have a favorite sport to watch?
Sports, Games It is good to spend time outdoors exercising.
@@ 3.3.1. Robot System
We utilized a Panda robot from Franka Emika. It is an industrial robotic arm with 7 degrees of freedom (7-DoF). The robot was restricted to its designated workspace and never intruded into the participant’s allocated area. Its movements were preprogrammed to execute specific sequences of actions. Among these, two involved pick-and-place motions: one to retrieve and deliver either a green or yellow PVC pipe to the participant, and the other to pick up a container holding a PVC pipe joint and subsequently move the container over a bin to discard the joint. The robot interacted with participants using text-to-speech communication. Its speech and physical actions were controlled by the experimenter via a Wizard-of-Oz (WoZ) approach (Riek, 2012).
@@ 3.3.2. Wizard-of-Oz Control
While monitoring the study interaction using a webcam, the experimenter used a remote graphical user interface (GUI) to click buttons that executed preprogrammed robot actions. For conversational speech, pre-written conversational phrases were assigned to specific alphanumeric codes. The experimenter could trigger one of the preprogrammed conversational speech phrases by inputting its code into a textbox in the GUI. Additionally, the experimenter could manually type text into the textbox to produce non-preprogrammed phrases via Google Text-to-Speech (TTS).
Figure 3. An overview of the collaborative tasks the participants completed with the robot. The team negative feedback was introduced halfway through the first and second task structures. To analyze changes in user behavior across instances of negative feedback, we delimit observation windows with respect to the negative feedback announcements. Window A is considered our baseline as this corresponds to the only portion of the interaction that has not yet been influenced by any negative feedback statements. In the lower half of the figure, we include an example of one user’s conversational interaction with the robot after the first negative feedback announcement (P10).
[…]
Top half of the figure contains an outline of the three tasks, with dots on a timeline signifying where a negative feedback announcement was given. The bottom of the figure contains images of the participant alongside their and the robot’s conversation as the task progresses. Next is the caption of the conversation, with S corresponding to the system announcement, R corresponding to the robot, and P corresponding to the participant; S: System announcement, team three, please work on improving you performance. P: Bring me a yellow pipe. R: Getting it now. P: Bring me a yellow pipe. R: Fetching it now. R: On another note, do you like to read in your spare time? (re-initiating small talk after two task requests) P: Sometimes but not for a while. I’ve been busy.